A conducting sphere of radius R and carrying a charge q is joined to another conducting sphere of radius 2R and carrying a charge -2q. The charge flowing between them will be


 


a) 1.5q


b) 0.67q


c) q


d) 1.33q

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

At equilibrium, the electrostatic potential on the two spheres is the same. Let Q be the charge on the smaller sphere.

Potential is directly proportional to charge/radius. thus, Q/R= (-q-Q)/2R

therefore, Q= -q/3, thus charge flow = q-(-q/3) =1.33q